# Equal Contribution   Equal Advising https://sirui-xu.github.io/InterDreamer/

### v1

@@ Abstract
Text-conditioned human motion generation has experienced significant advancements with diffusion models trained on extensive motion capture data and corresponding textual annotations. However, extending such success to 3D dynamic human-object interaction (HOI) generation faces notable challenges, primarily due to the lack of large-scale interaction data and comprehensive descriptions that align with these interactions. This paper takes the initiative and showcases the potential of generating human-object interactions without direct training on text-interaction pair data. Our key insight in achieving this is that interaction semantics and dynamics can be decoupled. Being unable to learn interaction semantics through supervised training, we instead leverage pre-trained large models, synergizing knowledge from a large language model and a text-to-motion model. While such knowledge offers high-level control over interaction semantics, it cannot grasp the intricacies of low-level interaction dynamics. To overcome this issue, we further introduce a world model designed to comprehend simple physics, modeling how human actions influence object motion. By integrating these components, our novel framework, InterDreamer, is able to generate text-aligned 3D HOI sequences in a zero-shot manner. We apply InterDreamer to the BEHAVE and CHAIRS datasets, and our comprehensive experimental analysis demonstrates its capability to generate realistic and coherent interaction sequences that seamlessly align with the text directives.
@@ 1 Introduction
Text-guided human motion generation [94] has made unprecedented progress through advancements in diffusion models [85, 86, 31], leading to synthesis outcomes that are more realistic, diverse, and controllable. This progress has further ignited increased interest in exploring expanded tasks related to text-guided human interaction generation, such as social interaction [54] and human-scene interaction [33]. However, many of these explorations are limited in that the dynamics of objects are not involved or cannot be controlled by text. Aiming to bridge such a gap, this paper undertakes the initiative to tackle a more challenging task – generating versatile 3D human-object interactions (HOIs) through language guidance, as illustrated in Fig. 1.
@@
Figure 1: InterDreamer can generate vivid 3D human-object interaction sequences guided by textual descriptions. Its zero-shot ability is achieved by integrating semantics and dynamics knowledge from large-scale text-motion data (upper left), a large language model (LLM) (bottom left), 3D human-object interaction database (upper middle), and interaction prior (bottom middle). We visualize the generated text-guided interaction sequence (upper right), with the beginning of the sequence unfolded (bottom right). More details are available in https://sirui-xu.github.io/InterDreamer/.
@@
While a direct solution, as suggested by concurrent work [71, 21, 50, 109], would be replicating the success observed in human motion generation and adopting a similar supervised approach for learning text-driven HOIs, it is not scalable. Indeed, even generating social or scene interactions heavily relies on extensive collections of text-interaction pair data [62, 25, 104, 54]. Scaling these methods to address the more complex HOIs outlined in our study could necessitate datasets of comparable magnitude. Achieving this goal appears unattainable by merely annotating existing 3D HOI datasets [7, 36, 34, 129, 22, 51, 136, 41], which are relatively limited in size. Although recent studies [71, 21, 50] have annotated some of these datasets, the volume of text-motion pairs still lags significantly behind that available for existing text-driven motion generation efforts.
@@
An intriguing question naturally emerges: what is the potential of zero-shot learning for text-conditioned HOI generation, which is the main focus of this paper. However, formulating the task in a zero-shot setting presents significant challenges, primarily due to the inability to directly learn the alignment between text and HOI dynamics. Our key observation then is that interaction semantics and dynamics can be decoupled. That is, the high-level semantics of an interaction, aligned with its textual description, can be informed by human motion and the initial object pose. Meanwhile, the low-level dynamics of the interaction – specifically, the subsequent behavior of the object – are governed by the forces exerted by the human, within the constraints of physical laws.
Motivated by these insights, we introduce InterDreamer – a novel framework that synergizes knowledge of interaction semantics and dynamics. As shown in Fig. 1, both of them do not necessarily need to be learned from text-interaction pairs, if they are decoupled, thus leading to the ability for zero-shot generation.
@@
The semantics of interaction, although not available through direct supervised training, can be harnessed from a variety of prior knowledge that is independent of text-interaction pair datasets. Specifically, to acquire semantically aligned human motion and initial interaction, we first consult a large language model (LLM), such as GPT-4 [67] and Llama 2 [95], to provide understandings including how humans typically use specific body parts in interactions with particular objects, by exploiting its in-context learning capability with few-shot prompting [10] and chain-of-thought prompting [107]. Both intermediate thoughts and the final thought are then used to (i) generate a semantically aligned human motion with a pre-trained text-to-motion model; and (ii) identify an initial object pose that is harmonious with the generated initial human pose and text description through a retrieval-based algorithm.
@@
While these large models can offer high-level motion semantics modeling, they lack crucial low-level dynamics knowledge. Nevertheless, by decoupling interaction dynamics from semantics, a key advantage emerges in our InterDreamer framework: interaction dynamics can be learned from motion capture data without the necessity of text annotations. We instantiate this idea by developing a novel world model, which predicts the subsequent state of an object affected by the interaction. The key here is to attain generalizable control signals. To do so, we exert control over the object through the motion of vertices on the human body. These vertices are solely sampled in regions where contact occurs, agnostic to the overall object shape and whole-body motion. Such abstraction empowers the model to learn the fundamental physics from a publicly available 3D HOI dataset [7]. The plausibility of generated interaction is further enhanced by a subsequent optimization procedure on synthesized human and object motion.
@@
To summarize, our contributions are: (i) We initiate a novel task of synthesizing whole-body interactions with dynamic objects guided by textual commands, without access to text-interaction pair data, which is the first to our knowledge. (ii) We introduce a novel framework that decomposes semantics and dynamics, and they can be integrated effortlessly. (iii) Our methodology harnesses knowledge from a large language model (LLM) and a text-to-motion model as external resources, alongside our proposed novel world model. Remarkably, the only component that requires training is the world model, underscoring the ease of use of our framework. Experimental results demonstrate that our zero-shot framework, InterDreamer, is capable of producing semantically aligned and realistic human-object interactions, and generalizes beyond existing HOI datasets.
@@ 2 Related Work
Text-Conditioned Human Motion Generation. Significant progress has been witnessed in human motion synthesis tasks, given different kinds of external conditions, including action categories [27, 73, 49, 2], past motions [125, 65, 5, 14, 117, 118], trajectories [39, 38, 80, 113, 97], scene context [11, 29, 99, 101, 100, 104, 33, 137, 138, 91, 132], and unconditional generation [76]. Recently, human motion synthesis guided by textual descriptions [75, 26, 74, 15, 130, 134, 128, 92, 1, 25, 42, 59, 77, 84, 20, 106, 135, 45, 123, 6, 141, 60] is popular and extended to various applications, including the text-conditioned generation of multiple-person [57, 105, 24] and human-scene interaction [33, 37, 17]. Our goal is to model human and object dynamics concurrently guided by text in a zero-shot setting.
@@
Human-Object Interaction Generation. Synthesizing hand-object interactions [53, 124, 139, 140, 126] and single-frame human-object interactions [112, 127, 102, 72, 32, 44] are popular topics and extended to zero-shot settings [52, 120, 40]. Recently, researchers explore whole-body dynamic interaction generation, in kinematic-based approaches [87, 88, 89, 110, 47, 133, 48, 119, 23, 18, 98, 79, 63, 64, 46, 116, 51, 109] and physics-based approaches [56, 13, 66, 30, 4, 121, 114, 115, 68, 8, 103, 19]. Current methods in HOI synthesis are often restricted by a narrow scope of actions, the use of non-dynamic objects, and a lack of comprehensive whole-body motion. Our work aims to generate diverse whole-body interactions with various objects, and enables control through language input. Recent datasets [90, 7, 36, 34, 129, 22, 51, 136, 41] provide the groundwork for research in this area, and concurrent efforts [71, 21, 50] demonstrate the feasibility of applying supervised learning methods through annotating datasets. However, the amount of data currently available falls short when compared to more extensive text-motion datasets [62, 25, 55]. This discrepancy in data volume limits the capability of supervised methods to capture the complexity of human-object interactions, motivating us to investigate the potential of zero-shot generation.
@@
External Knowledge from LLMs. Large language models (LLMs) are being used for advanced visual tasks, such as editing images based on instructions [9]. In digital humans, they are used to reconstruct 3D human-object interactions [102] and generate human motions [3, 122, 35, 134] as well as human-scene interactions [111]. Our approach is inspired by [102], which uses LLMs to infer contact body parts with a given object for reconstructing 3D human-object interactions – a task different from ours. Also, our approach utilizes more advanced LLMs, such as GPT-4 [67] or Llama 2 [95], not only to understand contact body parts but also to further narrow the distribution gap between the text descriptions and models in our pipeline. This is achieved by leveraging the in-context learning (ICL) of LLMs with the chain-of-thought prompting [107].
@@ 3 Methodology
Problem Formulation. Our goal is to synthesize a sequence of 3D human-object interactions $\boldsymbol{x}$ that corresponds to a descriptive text $p$. This sequence is a series of tuples $[(\boldsymbol{h}_{1},\boldsymbol{o}_{1}),(\boldsymbol{h}_{2},\boldsymbol{o}_{2}),\ldots,(\boldsymbol{h}_{M},\boldsymbol{o}_{M})]$, where $\boldsymbol{h}_{i}$ represents the human pose parameters defined by the SMPL model [58], and $\boldsymbol{o}_{i}$ defines the object pose in terms of its 3D spatial location and orientation. The sequence length $M$ is variable and is dynamically determined by our text-to-motion model based on the input text $p$. We do not require text-interaction paired data for training.
@@
Figure 2: An overview of our InterDreamer. (i) Our high-level planning analyzes the description using LLMs and provides guidance to the low-level control. (ii) Our low-level control includes a text-to-motion model that translates text into human actions $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{a}_{t+1}$, and an interaction retrieval model for extracting the object’s initial pose as the first state $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{s}_{1}$. (iii) Our world model executes the actions and outputs the next state $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{s}_{t+1}$ through dynamics modeling. An optimization process is coupled with the dynamics model, projecting the state and action onto valid counterparts $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{s}_{t+1}^{\ast}$ and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\boldsymbol{a}_{t+1}^{\ast}$. Solid arrows mean that the process is performed iteratively.
@@
Overview. Our framework, illustrated in Fig. 2, can be conceptualized as a Markov decision process (MDP). We begin by dividing the motion sequence into $T$ segments, each with $m$ frames, where $M=T\times m$. Object motion $\{\boldsymbol{o}_{i}\}_{i=1}^{M}$ is formulated as a sequence of environmental states $\{\boldsymbol{s}_{t}\}_{t=1}^{T}$, and human motion $\{\boldsymbol{h}_{i}\}_{i=1}^{M}$ is described as a sequence of actions $\{\boldsymbol{a}_{t}\}_{t=1}^{T}$ that interact with the environment. Under such an MDP setup, our pipeline starts with high-level planning $L$, which, given the textual interaction description $p$, deciphers the detailed context $g=L(p)$ (Sec. 3.1). Then, a text-to-motion model $\pi$ translates context $g$ into human actions iteratively, modeled as $\boldsymbol{a}_{t+1}\sim\pi(\boldsymbol{a}_{t+1}|\boldsymbol{s}_{t},\{\boldsymbol{a}_{i}\}_{i=1}^{t},g)$ (Sec. 3.2). An interaction retrieval model $R$ proposes an initial object state $\boldsymbol{s}_{1}$, based on the initial action $\boldsymbol{a}_{1}$ and context $g$ (Sec. 3.2). After that, a world model $P$ is trained to predict future states $\boldsymbol{s}_{t+1}$ from the current action and state (Sec. 3.3). Our world model further incorporates an optimization process with prior knowledge on human interactions encoded, for both state and action refinement (Sec. 3.4). Notably, the text-to-motion and world models are executed iteratively until text-to-motion generates an end frame.
@@ 3.1 High-Level Planning
Powered by LLMs’ strong reasoning capabilities as well as their common sense, our high-level planning $L$ yields interaction details $g=L(p)$ that cannot be directly extracted in textual descriptions $p$. The process undertaken by $L$ encompasses three steps: (i) Determining the object: The LLM is employed to translate the described objects into corresponding categories from a pre-defined list used in the world model. (ii) Determining initial human-object contact: The LLM infers the body parts involved in the interaction, drawing from a list defined in the SMPL model [58]. (iii) Reducing the distribution gap: The LLM bridges the distribution gap between the freeform textual input and the language used within the text-to-motion model training dataset [25]. This involves standardizing syntax and content according to specific guidelines we design.
In Fig. 2, we demonstrate the prompt used by the LLM, as well as examples, following the few-shot prompting [10]. We define the sequence of intermediate thoughts and the final thought, i.e., the answers to the three questions, as the detailed information $g=L(p)$, which guides the subsequent procedure.
Our high-level planning operates indirectly in the generation of interactions. Nonetheless, it plays a key role in effectively bridging the distribution gap between textual descriptions and subsequent interaction generation. In other words, it narrows the vast range of possible interactions into a more manageable distribution within the capabilities of our framework. As detailed in Sec. 4.1, our high-level planning incorporates GPT-4 [67] or Llama-2 [95] for evaluation.

### v2

# InterDreamer: Zero-Shot Text to 3D Dynamic Human-Object Interaction

@@ Abstract
Text-conditioned human motion generation has experienced significant advancements with diffusion models trained on extensive motion capture data and corresponding textual annotations. However, extending such success to 3D dynamic human-object interaction (HOI) generation faces notable challenges, primarily due to the lack of large-scale interaction data and comprehensive descriptions that align with these interactions. This paper takes the initiative and showcases the potential of generating human-object interactions without direct training on text-interaction pair data. Our key insight in achieving this is that interaction semantics and dynamics can be decoupled. Being unable to learn interaction semantics through supervised training, we instead leverage pre-trained large models, synergizing knowledge from a large language model and a text-to-motion model. While such knowledge offers high-level control over interaction semantics, it cannot grasp the intricacies of low-level interaction dynamics. To overcome this issue, we introduce a world model designed to comprehend simple physics, modeling how human actions influence object motion. By integrating these components, our novel framework, InterDreamer, is able to generate text-aligned 3D HOI sequences without relying on paired text-interaction data. We apply InterDreamer to the BEHAVE, OMOMO, and CHAIRS datasets, and our comprehensive experimental analysis demonstrates its capability to generate realistic and coherent interaction sequences that seamlessly align with the text directives.
@@ 1 Introduction
Text-guided human motion generation has made unprecedented progress through advancements in diffusion models [104, 105, 41, 131], leading to synthesis outcomes that are realistic, diverse, and controllable. This progress has ignited an increased interest in exploring expanded tasks related to text-guided human interaction generation, such as social interaction [69] and human-scene interaction [44]. However, many of these explorations are limited in that the dynamics of objects is not involved or text-guided. Aiming to bridge such a gap, this paper tackles a more challenging task – generating versatile 3D human-object interactions (HOIs) through language guidance, as illustrated in Figure 1.
@@
Figure 1: InterDreamer generates vivid 3D human-object interaction sequences guided by text descriptions, by synergizing semantics and dynamics knowledge from large-scale text-motion data (upper left), a large language model (bottom left), human-object interaction data (upper middle), and prior knowledge (bottom middle) from simple physics. We visualize the generated text-guided interaction sequence (upper right), with the beginning of the sequence unfolded (bottom right).
@@
Although a direct solution, as suggested by the concurrent work [90, 28, 65, 136, 140, 106], would be replicating the success observed in human motion generation and adopting a similar supervised approach for learning text-driven HOIs, it is not scalable. As can be observed, generating social or scene interactions is heavily dependent on extensive collections of text-interaction pair data [82, 34, 130, 69], and scaling these methods to address more complex HOIs outlined in our study could require datasets of comparable magnitude. Achieving this goal appears unattainable by merely annotating existing 3D HOI datasets [7, 47, 45, 169, 30, 66, 180, 53, 177, 170], which are relatively limited in size. Although recent studies [90, 28, 65, 155] have annotated some of these datasets, the volume of text-interaction pairs still lags behind that available for existing text-driven motion generation efforts.
@@
An intriguing question naturally arises: given the limited annotations of the text, what is the potential of learning for text-conditioned HOI generation without text supervision, which is the main focus of this paper. However, formulating the task in such a setting presents significant challenges, primarily due to the inability to directly learn the alignment between text and HOI dynamics. Our key observation is that interaction semantics and dynamics can be decoupled. That is, the high-level semantics of an interaction, aligned with its textual description, can be informed by human motion and the initial object pose. Meanwhile, the low-level dynamics of the interaction – specifically, the subsequent behavior of the object – is governed by the forces exerted by the human, within the constraints of physical laws. Motivated by these insights, we introduce InterDreamer – a novel framework that synergizes knowledge of interaction semantics and dynamics (Figure 1), both of which do not necessarily require learning from text-interaction pairs, if they are decoupled.
@@
The semantics of interaction, although not available through direct supervised training, can be harnessed from prior knowledge without text-interaction pair datasets. Specifically, to acquire semantically aligned interaction, we first consult a large language model (LLM), such as GPT-4 [87] and Llama 2 [119], to provide understanding including how humans typically use specific body parts in interactions with particular objects, by exploiting its in-context learning capability with few-shot prompting [10] and chain-of-thought prompting [134]. The intermediate thoughts and the final thought are then used to (i) generate semantically aligned human motion with a pre-trained text-to-motion model; and (ii) identify an initial object pose that is harmonious with the generated human pose and text description, following a philosophy similar to retrieval-augmented generation [62].
@@
While these large models can offer high-level motion semantic modeling, they lack crucial low-level dynamics knowledge. Nevertheless, by decoupling interaction dynamics from semantics, a key advantage emerges in our InterDreamer framework: interaction dynamics can be learned from motion capture data without the necessity of text annotations. We instantiate this idea by developing a world model, which predicts the subsequent state of an object affected by the interaction. The key here is to reach generalizable representations in different motion and objects. To do so, we exert control over the object through the motion of vertices on the human body. These vertices are solely sampled in regions where contact occurs, agnostic to the overall object shape and whole-body motion. Such abstraction empowers the model to learn the simple dynamics from a publicly available 3D HOI dataset BEHAVE [7], and generalize naturally to other datasets [66, 47]. The plausibility of the generated interaction is further enhanced by a subsequent optimization procedure on the synthesized human and object motion.
@@
To summarize, our contributions are: (i) We address the task of synthesizing whole-body interactions with dynamic objects guided by textual commands, achieving this notably without the need for paired text-interaction data, a novel paradigm to the best of our knowledge. (ii) We introduce a framework that decomposes semantics and dynamics, and they can be easily integrated. Specifically, it harnesses knowledge from a large language model (LLM) and a text-to-motion model as external resources, alongside our proposed world model. Remarkably, the only component that requires additional training is the world model, which highlights the ease of use of our framework. (iii) Experimental results demonstrate that our framework, InterDreamer, is capable of producing semantically aligned and realistic human-object interactions, and generalizes beyond existing HOI datasets.
@@ 2 Related Work
Text-Conditioned Human Motion Generation. Significant progress has been witnessed in human motion synthesis tasks, given different kinds of external conditions, including action categories [36, 92, 61, 2], past motion [163, 85, 5, 17, 149, 150, 109], trajectories [51, 50, 99, 145, 122, 31], scene context [12, 39, 125, 127, 126, 130, 44, 182, 183, 113, 175, 29, 112, 153], and without condition [95]. Recently, human motion synthesis guided by textual descriptions [93, 35, 94, 18, 172, 178, 168, 115, 1, 34, 54, 77, 96, 103, 24, 133, 179, 57, 160, 6, 187, 81, 185, 181, 49, 64, 26, 71, 162, 174] is popular and extended to various applications, including the text-conditioned generation of multiple-person [75, 132, 33, 63, 43] and human-scene interaction [44, 48, 21, 14]. Our goal is to model human and object dynamics concurrently guided by text.
@@
Human-Object Interaction Generation. Synthesizing hand-object interactions [68, 161, 184, 186, 165, 11, 167, 79, 80, 118, 137, 20, 73, 15] and single-frame human-object interactions [143, 166, 128, 91, 42, 56, 152, 154, 25] are popular topics and extended to zero-shot settings [67, 156, 52, 157]. Recently, researchers explore whole-body dynamic interaction generation, in kinematic-based approaches [107, 108, 110, 139, 59, 176, 60, 151, 32, 22, 123, 98, 83, 84, 58, 148, 66, 38, 136, 27, 140, 106] and physics-based approaches [72, 16, 86, 40, 4, 158, 146, 147, 88, 8, 129, 142, 23, 114, 74, 124, 78, 116]. Current methods in HOI synthesis are often restricted by a narrow scope of actions, the use of non-dynamic objects, and a lack of comprehensive whole-body motion. Our work aims to generate diverse whole-body interactions with various objects, and enables control through language input. Recent datasets [111, 7, 47, 45, 169, 30, 66, 180, 53, 138, 155, 164, 144, 170] provide the groundwork for research in this area, and concurrent efforts [90, 28, 65] demonstrate the feasibility of applying supervised learning methods via annotating datasets. However, the amount of data currently available fall short when compared to more extensive text-motion datasets [82, 34, 70]. This discrepancy in data volume limits the capability of supervised methods to capture the complexity of human-object interactions, motivating us to investigate the potential of zero-shot generation.
@@
External Knowledge from LLMs. Large language models (LLMs) are being used for advanced visual tasks, such as editing images based on instructions [9]. In digital humans, they are used to reconstruct 3D human-object interactions [128] and generate human motion [3, 159, 46, 178] as well as human-scene interactions [141]. Our approach is inspired by [128], which uses LLMs to infer contact body parts with a given object for reconstructing 3D human-object interactions – a task different from ours. Our approach utilizes GPT-4 [87] or Llama 2 [119], to not only understand contact body parts but also narrow the distribution gap between different tasks, and provide knowledge for interaction retrieval. This is accomplished by utilizing the in-context learning capabilities of LLMs [22] and their support for retrieval-augmented generation [62].
@@ 3 Methodology
Problem Formulation. Our goal is to synthesize a sequence of 3D human-object interactions $\bm{x}$ that satisfies a descriptive text $p$. This sequence is a series of tuples $(\bm{h}_{1},\bm{o}_{1}),(\bm{h}_{2},\bm{o}_{2}),\ldots,(\bm{h}_{M},\bm{o}_{M})]$, where $\bm{h}_{i}$ represents the human pose parameters defined in the SMPL model [76], while the shape of the human is unified the same as [34]. $\bm{o}_{i}$ defines the pose of the rigid object in terms of its 3D spatial position and orientation. The sequence length $M$ is variable and is dynamically determined by our text-to-motion model based on the input text $p$. We do not require text supervision for training.
@@
Figure 2: An overview of our InterDreamer. (i) Our high-level planning analyzes the description using LLMs and provides guidance to the low-level control. (ii) Our low-level control includes a text-to-motion model that translates text into human actions $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{a}_{t+1}$, and an interaction retrieval model that extracts the object’s first state $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{s}_{1}$. (iii) Our world model executes actions to output the next state $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{s}_{t+1}$. We achieve this by abstracting the problem as predicting the motion of contact vertices – represented by red spheres for humans and blue spheres for objects on the top right – using human vertices as controls for the prediction of object vertices. An optimization process is coupled with the dynamics model, projecting the state and action onto valid counterparts. Solid arrows mean that the process is performed iteratively.
@@
Overview. Our framework, illustrated in Figure 2, can be conceptualized as a Markov decision process (MDP). We begin by dividing the motion sequence into $T$ segments, each with $m$ frames, where $M=T\times m$. Object motion $\{\bm{o}_{i}\}_{i=1}^{M}$ can be seen as a sequence of environmental states $\{\bm{s}_{t}\}_{t=1}^{T}$, and human motion $\{\bm{h}_{i}\}_{i=1}^{M}$ is described as a sequence of actions $\{\bm{a}_{t}\}_{t=1}^{T}$ that interact with the environment. Under such an MDP setup, our framework starts with high-level planning $L$, which deciphers textual interaction description $p$ by $g=L(p)$ (Sec. 3.1). Then, a text-to-motion model $\pi$ translates context $g$ into human actions, modeled as $\bm{a}_{t+1}\sim\pi(\bm{a}_{t+1}|\bm{s}_{t},\{\bm{a}_{i}\}_{i=1}^{t},g)$ (Sec. 3.2). The interaction retrieval $R$ proposes an initial object state $\bm{s}_{1}\sim R(\bm{s}_{1}|\bm{a}_{1},g)$, based on the initial action $\bm{a}_{1}$ and context $g$ (Sec. 3.2). After that, a world model $P$ is trained to predict future states $\bm{s}_{t+1}\sim P(\bm{s}_{t+1}|\bm{a}_{t},\bm{s}_{t},\bm{a}_{t+1})$ from the current action and state (Sec. 3.3). Our world model incorporates an optimization process, for both state and action refinement (Sec. 3.4). Notably, the text-to-motion and world models are executed iteratively until text-to-motion generates an end frame.
@@ 3.1 High-Level Planning
Leveraging LLMs’ strong reasoning capabilities and inherent common sense, our high-level planning $L$ yields interaction details $g=L(p)$ that cannot be naïvely extracted in textual descriptions $p$. The process undertaken by $L$ encompasses three steps: (i) Determining the object: the LLM is employed to translate described objects into corresponding categories from a predefined list. (ii) Determining initial human-object contact: the LLM infers the body parts involved in the interaction, drawing from a list defined in the SMPL model [76]. And most importantly, (iii) reducing the distribution gap: the LLM bridges the distribution gap between the free-form textual input and the language used within the training data of the text-to-motion model [34]. This involves standardizing syntax and content according to designed guidelines. In Figure 2, we demonstrate the prompt we used with the few-shot prompting [10]. We define intermediate thoughts and the final thought, i.e., answers to three questions, as detailed information $g=L(p)$, which guides the subsequent procedure, structuring the entire framework with a philosophy similar to retrieval-augmented generation [62]. Our high-level planning operates indirectly in the generation of interactions. Nonetheless, it narrows the vast range of possible interactions in the real world into a more manageable distribution within the capabilities of our framework. We incorporate GPT-4 [87] and Llama-2 [119] for evaluation.